# Black-Box Access is Insufficient for Rigorous AI Audits

### v1

@@ 7. Discussion
However, absent legal requirements, developers have strong incentives to limit access granted to external auditors. Thus far, external audits of state-of-the-art AI systems, when they have occurred, have been black-box (to public knowledge) (METR, 2023; Kinniment et al., 2023; OpenAI, 2023b; Anthropic, 2023; Touvron et al., 2023), suggesting a failure of existing incentive structures to provide greater access to auditors. Developers are typically reluctant to provide more permissive access to their models and related resources (Raji et al., 2022). This may stem from concerns that information collected through white- and outside-the-box audits could be leaked (Bommasani et al., 2023) (though this is addressable–see Section 6). Furthermore, it could expose a system’s lackluster performance, vulnerabilities, or poor risk management processes by developers, which could lead to reputational harm and possibly legal liability (Lemley et al., 2023).
Current black-box audits may set a precedent for future norms. Established norms frequently become “sticky” and entrenched in regulatory regimes (Nielson, 2018). Accordingly, the current norm for black-box audits (METR, 2023; Kinniment et al., 2023; OpenAI, 2023b; Anthropic, 2023; Touvron et al., 2023) may set the future standard. Furthermore, in policy debates about audits, industry actors have also lobbied for limiting external auditors to black-box access (Google, 2021). Without sufficient access and resources, non-industry researchers will struggle to iterate upon methods for more thorough audits (Bucknall and Trager, 2023). Over time, this may limit or bias public understanding of AI systems. A lack of open research on transparency tools and the view that there is little social benefit from greater transparency are mutually reinforcing.
@@
Conclusion: We have argued that providing auditors with white-box and thorough outside-the-box access to systems is feasible and allows for more meaningful oversight from audits. We draw two conclusions. First transparency regarding model access and evaluation methods is necessary to properly interpret the results of an AI audit. Second, white- and outside-the-box access allow for substantially more scrutiny than black-box access alone. When higher levels of scrutiny are desired, audits should be conducted with higher levels of access. Finally, we emphasize that white-box and thorough outside-the-box access are necessary but not sufficient. Audits can and do fail for many reasons. Without careful institutional design, the incentives of developers and auditors may result in audits that do not consistently align with public interest (Raji et al., 2020b, 2022; Costanza-Chock et al., 2022; Anderljung et al., 2023b). In Appendix D, we examine additional ways in which the quality of audits can be compromised
@@ Acknowledgements.
Appendix A Goals for External Audits
Audits involve formally evaluating systems to assess risks, compliance with standards and regulations, and other desiderata of interest to stakeholders. High-quality audits from independent, external auditors can serve several purposes:
@@ Appendix A Goals for External Audits
  • Incentivizing responsible development: When individual components of the development process are insufficiently documented, information necessary to contextually assess risks is lost [Khlaaf, 2023]. Audits can assess the sufficiency of internal controls, risk assessment, and documentation [Raji et al., 2020b, Brundage et al., 2020, Koessler and Schuett, 2023, Schuett, 2023]. Greater accountability for internal practices incentivizes auditees to spend more effort on risk mitigation and documentation [Wagner and Dittmar, 2006], especially when facing penalties or public scrutiny [Goh and Li, 2013, Yew and Hadfield-Menell, 2022].
@@
  • Enabling fixes to technical problems: When problems are found during an audit, developers can then work to address them [Raji and Buolamwini, 2019]. External audits can also identify risk factors that might merit further guardrails on deployment, closer monitoring of deployed systems, or follow-up studies of user impacts.
  • Balancing transparency and security: Keeping systems entirely secret is maximally secure but prevents external scrutiny. Open-sourcing them allows for maximal scrutiny but can proliferate proprietary or misusable systems [Seger et al., 2023]. Audits offer a middle ground that allows for some transparency and independent risk assessment with high security.
@@
Appendix B Technical Assistance for Auditors
In Section 5, we discuss how outside-the-box access to information can help auditors conduct audits more effectively. However, for similar reasons, access to technical assistance can also be useful. For example, one resource that auditors will often need, especially for large language models, is to computing infrastructure [Anderljung et al., 2023b]. Further, additional technical assistance from the developers’ engineers may also help because they have unique practical knowledge of working effectively with their models. This may include assistance with fine-tuning, developing realistic test cases (e.g., [Zhou et al., 2023]), or integrating models with external tools that enhance capabilities to resemble real-world usage [Lee et al., 2023, Park et al., 2023b, Davidson et al., 2023, Naihin et al., 2023]. Past experience with AI audits has highlighted the value of technical assistance from developers.
@@ Appendix B Technical Assistance for Auditors
    –Anthropic on their audit by METR (formerly ARC Evals) [Anthropic, 2023]
@@
Appendix C Supporting Innovation on Auditing Tools
White-box tools for studying AI systems have long been a topic of technical interest, but research on methods often struggles to keep up with the scale and capabilities of AI systems [Bengio et al., 2023]. There are gaps between the capabilities of white-box evaluation tools and what auditors will need from them. More progress on both foundational research and practical tools will be useful, especially for state-of-the-art large language models because of their unique versatility and complexity.
Basic research: Current methods have provided useful insights. However, developing a detailed mechanistic understanding is not yet possible in state-of-the-art models. More progress in the basic science of neural networks and efforts to study their inner workings will help further research on evaluation techniques. This will require progress on both developing more intrinsically understandable systems and techniques to interpret trained ones [Carvalho et al., 2019, Räuker et al., 2023].
Practical tools: The goal of research on evaluation techniques is to produce methods that can be effectively used off-the-shelf by auditors. In the adversarial attack literature, benchmarks have largely focused on fooling networks with small perturbations to inputs instead of eliciting harm via more real-world features [Zhang et al., 2019, Roth et al., 2021, Hendrycks et al., 2021, Kaufmann et al., 2023]. In the interpretability literature, few benchmarks connected to practical tasks exist, with it being common to judge techniques based on researcher intuition [Dong et al., 2017, Miller, 2019, Krishnan, 2020, Adebayo et al., 2020, Räuker et al., 2023, Casper et al., 2023e]. Given the increasing scale and complexity of modern AI systems, developing more effective evaluation tools poses a challenge. Fortunately, open-source and API access to advanced AI systems has enabled progress on evaluation tools. However, no technique for benchmarking evaluation tools is more directly informative than applications on real systems. Partnerships between researchers and developers can facilitate these.
Secure auditing infrastructure: As discussed in Section 6, granting auditors white-box access to systems via application programming interfaces or secure research environments can reduce the risk of leaks. However, because norms for AI audits have not yet been established, there is little infrastructure for conducting audits securely. For example, efforts like the US National Deep Inference Facility project [United States National Science Foundation, 2023] could make more resources available to auditors. Establishing better tools and protocols is another priority [Lucaj et al., 2023]. At the same time, it will be key to establish norms and a regulatory framework around AI audits, as has been done in other industries with audits.
Appendix D Beyond Access: Other Needs for Rigorous Audits.
@@ Appendix C Supporting Innovation on Auditing Tools
Poorly-resourced audits: Working with state-of-the-art AI systems and effectively evaluating them requires compute and technical expertise. While developing and commercializing advanced AI systems can be lucrative, searching for problems with them might not be profitable or financially sustainable. Existing audits have largely relied on private funding (e.g., [METR, 2023, Kinniment et al., 2023, OpenAI, 2023a, Anthropic, 2023]), rather than public funding or other more sustainable, reliable, and diversified sources of funding.
@@
Narrowly-scoped audits: Audits may omit important evaluations. For example, early audits of GPT-4 have focused on risk-related capabilities [METR, 2023, Kinniment et al., 2023, OpenAI, 2023a] but did not appear to include external evaluation regarding other concerns such as robustness to adversarial attacks; potential for misinformation; demographic representation; or impacts on societal welfare, democracy, discrimination, and equality. Another way in which auditing can be narrow in scope is if it only occurs pre-deployment. A “black cloud” system with ever-changing components is even more difficult to evaluate than a black box.
@@
Conflicts of interest: Auditors may face pressure to refrain from insisting on sufficient access or conducting sufficiently rigorous audits. Auditor conflicts of interest, including collusion with auditees [White, 2010, Bolton et al., 2012] are well-known and long-standing problems [Goldman and Barlev, 1974, Moore et al., 2006]. They stem in part from the typical payment structure of auditors: auditors that produce more favorable evaluations – including due to receiving inadequate or incomplete information from audit targets – are often preferred over other auditors, leading companies to “opinion-shop” [Lennox, 2000] for comparatively lax evaluations. This can trigger a race to the bottom in which audits become progressively less rigorous and less informative [Baghai and Becker, 2020, Anderljung et al., 2023b]. This type of dynamic could emerge in the absence of adequate regulatory structures. For example, recent audits of state-of-the-art language models from OpenAI [OpenAI, 2023a] and Anthropic [Anthropic, 2023] were conducted on a voluntary basis by the Model Evaluation and Threat Research organization (METR, formerly named ARC Evals) [METR, 2023], which maintains a close relationship with both companies the details of which are not publicly disclosed.
@@
Cosmetic compliance: Absent clear legal requirements, companies have an incentive to prioritize cosmetic compliance with good practices [Krawiec, 2003], a form of cheap talk [Farrell and Rabin, 1996] or virtue signaling [Westra, 2021] in which audit targets create a superficial (yet misleading) appearance of good faith cooperation.
@@
Regulatory capture: While governance regimes that bolster auditing standards and procedures may appear promising, they, too, can be undermined. Studies in the field of organizational science demonstrate that companies respond strategically to interventions, employing a variety of operational, political, and legal tactics [Oliver, 1991] including supporting biased research [Abdalla and Abdalla, 2021]. In its simplest form, companies may selectively disclose audit-relevant information [Marquis et al., 2016, Hess, 2019], enabling them to game outcomes, including in AI audits [Raji et al., 2022]. Meanwhile, more sophisticated and well-resourced companies can shape the underlying audit criteria, metrics, and institutions, including by selecting which auditors have privileged access to information and which do not. Legal sociologists describe this symbiotic relationship between regulators and regulated entities as “legal endogeneity”: it is precisely the actors that law seeks to control that end up controlling the law [Edelman, 1992, 2016, Waldman, 2019, Abdalla and Abdalla, 2021]. AI audits are especially susceptible to these dynamics because the relevant standards are currently unclear [Costanza-Chock et al., 2022] and audit tools are bespoke and applied inconsistently across different developers and domains [Costanza-Chock et al., 2022].

### v2

@@ 7. Discussion
However, absent legal requirements, developers have strong incentives to limit access granted to external auditors. Thus far, external audits of state-of-the-art AI systems, when they have occurred, have been black-box (to public knowledge) (METR, 2023; Kinniment et al., 2023; OpenAI, 2023a; Anthropic, 2023; Touvron et al., 2023), suggesting a failure of existing incentive structures to provide greater access to auditors. Developers are typically reluctant to provide more permissive access to their models and related resources (Raji et al., 2022). This may stem from concerns that information collected through white- and outside-the-box audits could be leaked (Bommasani et al., 2023) (though this is addressable–see Section 6). Furthermore, it could expose a system’s lackluster performance, vulnerabilities, or poor risk management processes by developers, which could lead to reputational harm and possibly legal liability (Lemley et al., 2023).
Current black-box audits may set a precedent for future norms. Established norms frequently become “sticky” and entrenched in regulatory regimes (Nielson, 2018). Accordingly, the current norm for black-box audits (METR, 2023; Kinniment et al., 2023; OpenAI, 2023a; Anthropic, 2023; Touvron et al., 2023) may set the future standard. Furthermore, in policy debates about audits, industry actors have also lobbied for limiting external auditors to black-box access (Google, 2021). Without sufficient access and resources, non-industry researchers will struggle to iterate upon methods for more thorough audits (Bucknall and Trager, 2023). Over time, this could limit or bias public understanding of AI systems. A lack of open research on transparency tools and the view that there is little social benefit from greater transparency are mutually reinforcing. Appendix C expands on how investment, research, and development into auditing techniques and infrastructure can facilitate further progress.
@@
Conclusion: We have argued that providing auditors with white-box and thorough outside-the-box access to systems is feasible and allows for more meaningful oversight from audits. We draw two conclusions. First transparency regarding model access and evaluation methods is necessary to properly interpret the results of an AI audit. Second, white- and outside-the-box access allow for substantially more scrutiny than black-box access alone. When higher levels of scrutiny are desired, audits should be conducted with higher levels of access. Finally, we emphasize that white-box and thorough outside-the-box access are necessary but not sufficient for rigor. Audits can and do fail for many reasons. Without careful institutional design, the incentives of developers and auditors may result in audits that do not consistently align with public interest (Raji et al., 2020b, 2022; Costanza-Chock et al., 2022; Anderljung et al., 2023b; Birhane et al., 2024; Ojewale et al., 2024). In Appendix D, we examine additional ways in which the quality of audits can be compromised.
Ethics Statement
Given the role of rigorous audits in improving accountability and representing public interests, we expect the foreseeable impacts of this paper to be positive. However, audits can fail to benefit the public for a variety of reasons. We discuss these challenges in Appendix D.
@@ Acknowledgements.
Appendix A Motivations for External Audits
Audits involve formally evaluating systems to assess risks, compliance with standards and regulations, and other desiderata of interest to stakeholders. High-quality audits from independent, external auditors have been motivated in multiple ways (Sharkey et al., 2024):
@@ Appendix A Motivations for External Audits
  • Incentivizing responsible development: When individual components of the development process are insufficiently documented, information necessary to contextually assess risks is lost (Khlaaf, 2023). Audits can assess the sufficiency of internal controls, risk assessment, and documentation (Raji et al., 2020b; Brundage et al., 2020; Koessler and Schuett, 2023; Schuett, 2023). Greater accountability for internal practices incentivizes auditees to spend more effort on risk mitigation and documentation (Wagner and Dittmar, 2006), especially when facing penalties or public scrutiny (Goh and Li, 2013; Yew and Hadfield-Menell, 2022).
@@
  • Enabling fixes to technical problems: When problems are found during an audit, developers can then work to address them (Raji and Buolamwini, 2019). External audits can also identify risk factors that might merit further guardrails on deployment, closer monitoring of deployed systems, or follow-up studies of user impacts.
  • Balancing transparency and security: Keeping systems entirely secret is maximally secure but prevents external scrutiny. Open-sourcing them allows for maximal scrutiny but can proliferate proprietary or misusable systems (Seger et al., 2023). Audits offer a middle ground that allows for some transparency and independent risk assessment with high security.
@@
Appendix B Technical Assistance as a form of Outside-the-Box Access
In Section 5, we discuss how outside-the-box access to information can help auditors conduct audits more effectively. However, for similar reasons, access to technical assistance can also be useful. For example, one resource that auditors will often need, especially for large language models, is computing infrastructure (Anderljung et al., 2023b; Ojewale et al., 2024). Further, additional technical assistance from the developers’ engineers may also help because they have unique practical knowledge of working effectively with their models. This may include assistance with fine-tuning, developing realistic test cases (e.g., (Zhou et al., 2023)), or integrating models with external tools that enhance capabilities to resemble real-world usage (Lee et al., 2023; Park et al., 2023b; Davidson et al., 2023; Naihin et al., 2023). Past experience with AI audits has highlighted the value of technical assistance from developers.
@@ Appendix B Technical Assistance as a form of Outside-the-Box Access
–Anthropic on their audit by METR (formerly ARC Evals) (Anthropic, 2023)
@@
Appendix C Innovation on Auditing Tools
White-box tools for studying AI systems have long been a topic of technical interest, but research on methods often struggles to keep up with the scale and capabilities of leading AI systems (Bengio et al., 2023). There are gaps between the capabilities of white-box evaluation tools and what auditors may need from them. More progress on both foundational research and practical tools will be useful for auditors, especially for state-of-the-art large language models because of their unique versatility and complexity.
Basic research: Current methods have provided useful insights. However, developing a detailed mechanistic understanding is not yet possible in state-of-the-art models. More progress in the basic science of neural networks and efforts to study their inner workings will help further research on evaluation techniques. This will require progress on both developing more intrinsically understandable systems and techniques to interpret trained ones (Carvalho et al., 2019; Räuker et al., 2023).
Practical tools: The goal of research on evaluation techniques is to produce methods that can be effectively used off-the-shelf by auditors. In the adversarial attack literature, benchmarks have largely focused on fooling networks with small perturbations to inputs instead of eliciting harm via more real-world features (Zhang et al., 2019; Roth et al., 2021; Hendrycks et al., 2021; Kaufmann et al., 2023). In the interpretability literature, few benchmarks connected to practical tasks exist, with it being common to judge techniques based on researcher intuition (Dong et al., 2017; Miller, 2019; Krishnan, 2020; Adebayo et al., 2020; Räuker et al., 2023; Casper et al., 2023d). Given the increasing scale and complexity of modern AI systems, developing more effective evaluation tools poses a challenge. Fortunately, open-source and API access to advanced AI systems has enabled progress on evaluation tools. However, no technique for benchmarking evaluation tools is more directly informative than applications on real systems. Partnerships between researchers and developers can facilitate these.
Secure auditing infrastructure: As discussed in Section 6, granting auditors white-box access to systems via application programming interfaces or secure research environments can reduce the risk of leaks. However, because norms for AI audits have not yet been established, there is little infrastructure for conducting audits securely. For example, efforts like the US National Deep Inference Facility project (United States National Science Foundation, 2023) could make more resources available to auditors. Establishing better tools and protocols is another priority (Lucaj et al., 2023). At the same time, it will be key to establish norms and a regulatory framework around AI audits, as has been done in other industries with audits.
Appendix D Beyond Access: Other Aspects of Rigorous Audits.
@@ Appendix C Innovation on Auditing Tools
Poorly-resourced audits: Working with state-of-the-art AI systems and effectively evaluating them requires compute and technical expertise. While developing and commercializing advanced AI systems can be lucrative, searching for problems with them might not be profitable or financially sustainable. Existing audits have largely relied on private funding (e.g., (METR, 2023; Kinniment et al., 2023; OpenAI, 2023b; Anthropic, 2023)), rather than public funding or other more sustainable, reliable, and diversified sources of funding.
@@
Narrowly-scoped audits: Audits may omit important evaluations. For example, early audits of GPT-4 have focused on risk-related capabilities (METR, 2023; Kinniment et al., 2023; OpenAI, 2023b) but did not appear to include external evaluation regarding other concerns such as robustness to adversarial attacks; potential for misinformation; demographic representation; or impacts on societal welfare, democracy, discrimination, and equality. Another way in which auditing can be narrow in scope is if it only occurs pre-deployment. A “black cloud” system with ever-changing components is even more difficult to evaluate than a black box.
@@
Conflicts of interest: Auditors may face pressure to refrain from insisting on sufficient access or conducting sufficiently rigorous audits. Auditor conflicts of interest, including collusion with auditees (White, 2010; Bolton et al., 2012) are well-known and long-standing problems (Goldman and Barlev, 1974; Moore et al., 2006). They stem in part from the typical payment structure of auditors: auditors that produce more favorable evaluations – including due to receiving inadequate or incomplete information from audit targets – are often preferred over other auditors, leading companies to “opinion-shop” (Lennox, 2000) for comparatively lax evaluations. This can trigger a race to the bottom in which audits become progressively less rigorous and less informative (Baghai and Becker, 2020; Anderljung et al., 2023b). This type of dynamic could emerge in the absence of adequate regulatory structures. For example, recent audits of state-of-the-art language models from OpenAI (OpenAI, 2023b) and Anthropic (Anthropic, 2023) were conducted on a voluntary basis by the Model Evaluation and Threat Research organization (METR, formerly named ARC Evals) (METR, 2023), which maintains a close relationship with both companies the details of which are not publicly disclosed.
Exclusion of under-represented viewpoints: Not all people agree on what behaviors from AI systems are harmful. As a result, audits can exclude under-represented groups if they are designed in a way that fails to take a wide range of interests into account (Birhane et al., 2022; Lazar and Nelson, 2023; Feffer et al., 2024). Improving meaningful participation (Shur-Ofry, 2023; Sorensen et al., 2024) and dialogue (Dobbe et al., 2019) among diverse stakeholders plays an indispensable role in improving fairness and representation.
@@
Cosmetic compliance: Absent clear legal requirements, companies have an incentive to prioritize cosmetic compliance with good practices (Krawiec, 2003), a form of cheap talk (Farrell and Rabin, 1996) or virtue signaling (Westra, 2021) in which audit targets create a superficial (yet misleading) appearance of good faith cooperation.
@@
Regulatory capture: While governance regimes that bolster auditing standards and procedures may appear promising, they, too, can be undermined. Studies in the field of organizational science demonstrate that companies respond strategically to interventions, employing a variety of operational, political, and legal tactics (Oliver, 1991) including supporting biased research (Abdalla and Abdalla, 2021). In its simplest form, companies may selectively disclose audit-relevant information (Marquis et al., 2016; Hess, 2019), enabling them to game outcomes, including in AI audits (Raji et al., 2022; Birhane et al., 2024). Meanwhile, more sophisticated and well-resourced companies can shape the underlying audit criteria, metrics, and institutions, including by selecting which auditors have privileged access to information and which do not. Legal sociologists describe this symbiotic relationship between regulators and regulated entities as “legal endogeneity”: it is precisely the actors that law seeks to control that end up controlling the law (Edelman, 1992, 2016; Waldman, 2019; Abdalla and Abdalla, 2021). AI audits are especially susceptible to these dynamics because the relevant standards are currently unclear (Costanza-Chock et al., 2022) and audit tools are bespoke and applied inconsistently across different developers and domains (Costanza-Chock et al., 2022).